A critical reaction in the production of energy to do work or drive chemical reactions in biological systems is the hydrolysis of adenosine triphosphate, ATP, to adenosine diphosphate, ADP, as described by

ATP(aq)+ H2O(l) \rightarrow ADP(aq) + H2PO4(aq)

for which ΔG°rxn = –30.5 kJ/mol at 37.0 °C and pH 7.0. Calculate the value of ΔGrxn in a biological cell in which [ATP] = 5.0 mM, [ADP] = 0.80 mM, and [HPO42–] = 5.0 mM.

\DeltaGrxn = ___ KJ/mol

Is the hydrolysis of ATP spontaneous under these conditions?

yes or no?

Please complete all of the question, not half
Answer #1

The reaction is

AT P(aq) + H2O(l) + ADP(aq) + H2PO4(aq)

The formula for Gm is,

AGrrn = AGº – RT In

Here, эр is given as -30.5 kJ/mol and is Gibbs free energy at unit concentrations. Q is the reaction quotient and is defined as ratio of concentrations of the products divided by that of reactants. T is temperature and R is the universal gas constant.

Converting эр it to J/mol, we will have,

AG = (-30.5kJ/mol) x (1000J/kJ) = -30500J/mol

AT P = 5.0m M = 5.0 x 10-3M

ADP = 0.80m M = 0.80 x 10-3M

H2PO4) = 5.0m M = 5.0 x 10-3M

Temperature, T is

T = 37°C = (37 + 273) K = 310K

R= 8.314J/mol K

ADP H2PO4 [ATP

Thus, from (1), we have,

AGrrn = AG – RT In A [ADP][H2P04) [ATP]

Putting all the values, we have,

(0.80 x 10-3)(5.0 x 10-3) AGrin = -30500J/mol-(8.314J/mol K)(310K) in (5.0 x 10-3) → AGrm = -30500J/mol - (2577.34J/mol) In (

Note that in the above formula, the concentrations have been put in non-dimensional form by dividing by 1 M

Thus, the value of Gm is -12.121 kJ/mol

Further, the hydrolysis will be spontaneous because Arm < 0